Solve the following word problem using any method (if the problem has two unknowns, it should be two answers):
1. Steve has 35 coins consisting of dimes and nickels. If the value of the coins is $3.20, then how many of each type does he have?
Determine the equation of each line (write in slope-intercept form):
1.The line through (0, 6) that is parallel to the line 3x + 5y = 15
2.The line through (8, 0) that is perpendicular to the line x + y = 3
Write equation in slope-intercept form
1. 10 - x - 2y = 0

Answer by venugopalramana(3286)

1. Steve has 35 coins consisting of dimes and nickels.
LET NICKELS=N...THEN DIMES ARE...35-N
THEIR VALUE...5*N+10*(35-N)=5N+350-10N=350-5N
If the value of the coins is $3.20,
350-5N=320
5N=350-320=30
N=30/5=6
SO DIMES ARE 35-6=29 AND NICKLS=6
then how many of each type does he have?
Determine the equation of each line (write in slope-intercept form):
1.The line through (0, 6) that is parallel to the line 3x + 5y = 15
PARALLEL LINE WILL BE 3X+5Y=K..IT GOES THOUGH (0,6)
3*0+5*6=K=30....
SO QN.IS 3X+5Y=30
2.The line through (8, 0) that is perpendicular to the line x + y = 3
PERPENDICULAR LINE WILL BE X-Y=K...IT GOES THROUGH.(8,0)....
8-0=K....HENCE EQN. IS..
X-Y=8
Write equation in slope-intercept form
1. 10 - x - 2y = 0.......
2Y=-X+10
Y=-X/2+5